OpenStudy (anonymous):

Hiiii guys :) I need help again with my math. Okay here's the question: Suppose you invest $950 at an annual interest rate of 6.5% compounded continuously. How much will you have in the account after 10 years? Round the solution to the nearest dollar. I know how to start the equation but I always get the answer wrong no matter what. D: The formula is Pe^rt.

OpenStudy (anonymous):

A=950(1.91555) A=$1820

OpenStudy (jdoe0001):

\(\bf B = Pe^{rt}\\ \quad \\ \textit{keep in mind that "r" or rate is in decimal format, that is }6.5\%=0.065\\ \quad \\ B = 950\cdot e^{0.065\cdot 10}\implies B= 950\cdot e^{0.65}\)